# A switching identity for cable-graph loop soups and Gaussian free fields

### v2

@@ 1. Introduction
The paper [57] outlined some results and conjectures about the geometry of large sign clusters in ${\mathbb{Z}}^{d}$, and how these features depend on the spatial dimension $d$. In this direction, it is worth mentioning that intuitively speaking, when the dimension $d$ is large, the longer loops in a Brownian loop-soup become much less frequent, so that one expects the picture to resemble more that of ordinary critical percolation in high dimensions, whereas when $d\leq 5$, the large loops will play an important role in the appearance of large clusters, so that one is in a very different universality class than critical Bernoulli percolation. Indeed, when $d=2$, it has been shown that the loop-soup clusters (and sign-clusters) can be described via CLE4 loop ensembles (and their variants) in the scaling limit [33, 4]. There has been a fairly intense recent activity in the study of the large-scale properties (and exponents) of the sign clusters when $d\geq 3$ (see for instance [37, 13, 43, 11] – and in 2024 alone, it has been the main focus of [7, 8, 9, 20, 13, 14]).
@@
To our knowledge, this is essentially the only case (except for tree-like graphs) where a percolation measure conditioned on the existence of a connection has such a simple description. It is a rather powerful and useful tool: It allows to revisit/simplify the proofs of some of the existing results on cable-graph GFFs and their geometry, and gives access to further new results. Some of its immediate consequences that are given in Section 4 deal with the case where one of the points is sent to infinity. Let us state just one of these in this introduction namely the existence and a simple description of the incipient infinite cluster (IIC) for this percolation of loops model and leave further comments and results in Sections 4 and 6. We consider here $G$ to be ${\mathbb{Z}}^{d}$ for $d\geq 3$ but as we will discuss in Section 4.3, this result is in fact valid for a wide class of infinite transient graphs. We will use here the probability measure on “excursions from the origin to infinity” (which can be obtained by taking an unconditioned Brownian motion starting from $x$, and keeping only its part after its last visit of the origin:
@@
Let us conclude this introduction with some informal comments about why such a striking result has not been uncovered before. Indeed, once one guesses that the statement holds, it is not really difficult to prove! The main reason is arguably simply that the bijection between configurations with even and with odd number of excursions is not trivial (which provides an excuse to the community – or at least to me – to why it was not so easy to guess the result), especially when one looks at it from the lens of Dynkin’s isomorphism theorem without its loop-soup interpretation. Also, as we have already mentioned, it is not an observation that comes from discrete models – ideas from the continuum are needed. We note that it is not the first time that work on the (conformally invariant) two-dimensional continuous settings did lead mathematicians on the trail of general basic results about the GFF and loop-soups on cable graphs: The loop-soup itself was first introduced in the continuum setting in [27] in relation to the scaling limit of loop-erased random walk and the conformal restriction properties developed in [25], and its relation to the continuum two-dimensional GFF (via the combination of [49, 40] and [50]) was derived before Le Jan [29, 30] pointed out the direct relation between the square of the discrete GFF and the discrete loop-soups in general transient graphs or the direct basic relation with the loops erased in Wilson’s algorithm was worked out (see e.g. [24, 60]). This also holds in the present case: In [28], we were zooming in on some decompositions of two-dimensional loop-soup clusters given part of their boundaries (following earlier considerations in [46, 44, 45]), where some quizzing parity issues naturally popped up. A similar-looking switching identity appeared in [28] [in the latter section of that paper, we derive an identity in the (more involved) continuous two-dimensional case and for rectangular domains with some very special boundary conditions, so the points $x$ and $y$ in some sense correspond to two opposite vertical sides of the rectangle and the excursions from $x$ to $y$ do correspond to horizontal crossings]. This in turn led to the trail of the type of bijection that we will describe in Section 5 and also to the realisation that this type of switching result was actually working already at the simpler level of cable-graph GFFs and not only for some specific values of the values $a$ and $b$. So, again a somewhat convoluted route – with a special role played by the considerations, arguments and questions raised in [46, 28] coming from the continuum two-dimensional world.
@@ 3.2. Relation with the statements in [28]
3.3. The cable-graph ${\mathbb{R}}_{+}$
Let us make some remarks about the case where the cable-graph is ${\mathbb{R}}_{+}$. In this case, we in fact discussing squares of Bessel processes of integer dimensions in connection with Ray-Knight Theorem type identities. One indeed ends up with results in the spirit of Marc Yor’s beloved identities in law for excursions, meanders etc. as described in [61] (and it is nice to see the probabilistic worlds of Aizenman and Yor merging in this way!).
@@ 3.3. The cable-graph ${\mathbb{R}}_{+}$
  • So, nothing new so far, but the statement for non-negative $a$ and $b$’s provides some further “Pitman-Williams-Yor-type” identities in law involving Bessel bridges with randomly chosen dimensions in $\{3,7,11\ldots\}$. More precisely, if we now consider the cable-graph $[z,\infty)$ (for some $z<0$), $x=0$ and $y=1$, we obtain that for some explicit constant $C$:
@@
    Marc Yor would surely have come up instantaneously with proofs along the lines that we give in this section via explicit Laplace transform computations. It is interesting to see how his favourite techniques can be also reinterpreted via loop-soup decompositions. This is the type of “quest for mysteriously hidden pathwise explanations” of identities in law between functionals of stochastic processes that he was fond of.
@@ 5. Outlook: An explicit bijection via “peeling”
This question is reminiscent of papers by Warren, Yor, Aldous and others [54, 55, 3] that were studying the conditional distribution of one Brownian motion (or excursion) up to some given (stopping) time $T$ given its local time profile at $T$ – one idea being to try to progressively draw part of the Brownian motion, that then eats up progressively the allotted local time, and to describe the joint dynamics of the pair (Brownian motion, remaining allotted local time). Here, the question is how to do something similar for collections of Brownian loops.
@@
The obtained trajectory is then the “reversed vertex reinforced jump-process” as introduced by Sabot and Tarrès in [48] and further studied in [34]. In the limit where one lets $K\to\infty$, it is explained by Lupu, Sabot and Tarrès [35] that the process converges to a particular continuous self-interacting process. This process is absolutely continuous with respect to Brownian motion i.e., it is a Brownian motion with a rather complicated drift that is determined by its “remaining local time function” (the initial local time profile minus the local time of the past of the process). This updated remain profile will decrease when the process goes along. This process is well-defined and the convergence to this continuous process holds true up to the possible first time at which the local time at the current position of the process hits 00. Note that after that time, the process can not come back to this point anymore (as it would otherwise create a positive local time at this point, which is not possible), so that it has “to decide” in which direction to continue. Since this decision is “locally made” based on the local time profile, it is quite natural to guess that it will choose one of the two directions with equal probability.
@@ 6. Some first consequences for $n$-point functions and other features
One general remark is that up to the present paper, the study of large sign clusters in ${\mathbb{Z}}^{d}$ was focusing first on the derivation of connection probabilities between points, or between points and sets in the case of one-arm probabilities. For instance, a large part of the recent papers [13, 14, 15, 11, 7, 8, 9] is devoted on establishing bounds for the “one-arm” estimates, i.e. for the behaviour of the probability of the event $A_{N}$ that the sign cluster containing a given point is not contained in the ball of radius $N$ around this point when $N\to\infty$). Those proofs were mostly building on Lupu’s two-point function (and in some case on an extension from [37]) on the one hand, and on insightful delicate explorations of the GFF, involving very non-trivial renormalization arguments. With the multiple-point function estimates in hand (that they then subsequently obtain), it is then possible (see in particular very recent papers [15, 9]) to extract information about the large-scale properties of “typical” large clusters. This last step (from $n$-point functions to the properties of clusters) is not really specific to this loop-percolation case (see for instance the considerations of [2] for Bernoulli percolation).
@@ 6.3. The continuum versions
We conclude with the following point: In dimensions $d=3,4,5$, one expects (see [57]) the collection of loop-soup clusters to have a scaling limit when the mesh of the lattice goes to 00 (just as in $d=2$, see [33]), that will be coupled with a realization of the Brownian loop-soup in the continuum [and the properties of this scaling limit are then directly related with the large-scale properties of the models on cable-graphs]. In that context, given that the mass of excursions joining two far-away points in the cable-graph goes to 00 in the scaling limit, one expects that in the continuum, the switching property would hold in a similar way to Theorem 1: Conditioning two points $x$ and $y$ to be in the same “continuum sign-cluster” would amount to just adding one Brownian excursion joining these two points. This will already be interesting in dimension 2 (and a bounded domain). We plan to address such “switching property in the continuum setting results” in subsequent work [59].
@@ Acknowledgments.
This research has been supported by a Research Professorship of the Royal Society. I thank Juhan Aru and Hugo Duminil-Copin for their comments on a very preliminary version of this paper.

### v3

@@ 1. Introduction
The paper [63] outlined some results and conjectures about the geometry of large sign clusters in ${\mathbb{Z}}^{d}$, and how these features depend on the spatial dimension $d$. In this direction, it is worth mentioning that intuitively speaking, when the dimension $d$ is large, the longer loops in a Brownian loop-soup become much less frequent, so that one expects the picture to resemble more that of ordinary critical percolation in high dimensions, whereas when $d\leq 5$, the large loops will play an important role in the appearance of large clusters, so that one is in a very different universality class than critical Bernoulli percolation. Indeed, when $d=2$, it has been shown that the loop-soup clusters (and sign-clusters) can be described via CLE4 loop ensembles (and their variants) in the scaling limit [37, 6]. There has been a fairly intense recent activity in the study of the large-scale properties (and exponents) of the sign clusters when $d\geq 3$ (see for instance [41, 15, 48, 13] – and in 2024 alone, it has been the main focus of [9, 10, 11, 23, 15, 16, 22]).
@@
To our knowledge, this is essentially the only case (except for tree-like graphs) where a percolation measure conditioned on the existence of a connection has such a simple description. It is a rather powerful and useful tool: It allows to revisit/simplify the proofs of some of the existing results on cable-graph GFFs and their geometry, and gives access to further new results. Some of its immediate consequences that are given in Section 4 deal with the case where one of the points is sent to infinity. Let us state just one of these in this introduction namely the existence and a simple description of the incipient infinite cluster (IIC) for this percolation of loops model and leave further comments and results in Sections 4 and 6. We consider here $G$ to be ${\mathbb{Z}}^{d}$ for $d\geq 3$ but as we will discuss in Section 4.3, this result is in fact valid for a wide class of infinite transient graphs. We will use here the probability measure on “excursions from the origin to infinity” (which can be obtained by taking an unconditioned Brownian motion starting from $x$, and keeping only its part after its last visit of the origin):
@@
  5. (5)
    One can a posteriori detect aspects of this general switching property in a number of recent works on the GFF, especially those with the loup-soup perspective. This includes in the continuous and conformally invariant two-dimensional setting the papers [51, 32] that we will further comment on, as well as [25] [Remark 1.12 in that paper can be viewed as the scaling limit i.e., as the continuum version of Theorem 1 in this two-dimensional setting]. In the cable-graph setting, as we shall discuss in Section 3.3, an identity in law about bridges of three-dimensional Bessel processes by Pitman and Yor from the 1982 paper [47] can actually be interpreted as the switching property for the cable graph consisting of just one edge. Another relevant example is the paper [1] (see in particular Theorem 1.4 there that one could then use by adding an artificial additional edge bewtween two points to a cable graph) that also builds on the considerations [62, 40] (as part of the present paper does) that describes the number of excursions along one edge in the cable graph, where the conditioned Poisson random variables appear.
  6. (6)
    Both the GFF and its square (via the loop-soup representation) do satisfy a spatial Markov property. The switching property in some sense allows to reconcile the apparent parity-type contradiction that the combination of these two Markov properties do seem to create. Of course, one can also view it the other way around (and this is one way the proofs will go), namely that the switching property is a consequence of the combination of these two Markov properties.
@@
Let us conclude this introduction with some informal comments about why such a striking general result has not been uncovered before. Indeed, once one guesses that the statement holds, it is not really difficult to prove (and to come up with different proofs)! And as already mentioned, one can a posteriori see it “not far from the surface” at various places in the literature. The main reason is arguably simply that the bijection between configurations with even and with odd number of excursions is not trivial (which provides an excuse to the community – or at least to me – to why it was not so easy to guess the result), especially when one looks at it from the lens of Dynkin’s isomorphism theorem without its loop-soup interpretation. Also, as we have already mentioned, it is not an observation that comes from discrete models – ideas from the continuum are needed. We note that it is not the first time that work on the (conformally invariant) two-dimensional continuous settings did lead mathematicians on the trail of general basic results about the GFF and loop-soups on cable graphs: The loop-soup itself was first introduced in the continuum setting in [31] in relation to the scaling limit of loop-erased random walk and the conformal restriction properties developed in [29], and its relation to the continuum two-dimensional GFF (via the combination of [54, 44] and [55]) was derived before Le Jan [33, 34] pointed out the direct relation between the square of the discrete GFF and the discrete loop-soups in general transient graphs or the direct basic relation with the loops erased in Wilson’s algorithm was worked out (see e.g. [28, 66]). This also holds in the present case: In [32], we were zooming in on some decompositions of two-dimensional loop-soup clusters given part of their boundaries (following earlier considerations in [51, 49, 50]), where some quizzing parity issues naturally popped up. A similar-looking switching identity appeared in [32] [in the latter section of that paper, we derive an identity in the (more involved) continuous two-dimensional case and for rectangular domains with some very special boundary conditions, so the points $x$ and $y$ in some sense correspond to two opposite vertical sides of the rectangle and the excursions from $x$ to $y$ do correspond to horizontal crossings]. This in turn led to the trail of the type of bijection that we will describe in Section 5 and also to the realisation that this type of switching result was actually working already at the simpler level of cable-graph GFFs and not only for some specific values of the values $a$ and $b$. So, again a somewhat convoluted route – with a special role played by the considerations, arguments and questions raised in [51, 32] coming from the continuum two-dimensional world.
@@ 2. Proof of the switching property
Remark 9.
The paper [1] by Elie Aidékon contains some closely related considerations, see in particular his Theorem 1.4 where parity shows up – it does not address the switching property in terms of conditioning on long connections, but one a posteriori detect the switching property hiding near-by – one can for instance try to see what happens when one adds an additional edge to his setting. As we shall see, he also looked into the direction that we will describe in Section 5.
@@ 3.2. Relation with the statements in [32]
3.3. The cable-graphs ${\mathbb{R}}_{+}$ and $[0,1]$
Let us make some remarks about the case where the cable-graph is ${\mathbb{R}}_{+}$. In this case, we in fact discussing squares of Bessel processes of integer dimensions in connection with Ray-Knight Theorem type identities, and the parity and switching properties become reinterpretations of some of Marc Yor’s beloved identities in law for excursions, meanders etc. of the type described in [67] (and it is nice to see the probabilistic worlds of Aizenman and Pitman-Yor merging in this way!).
@@ 3.3. The cable-graphs ${\mathbb{R}}_{+}$ and $[0,1]$
  • The statement for non-negative $a$ and $b$’s provides some further “Pitman-Williams-Yor” identities in law involving Bessel bridges with randomly chosen dimensions in $\{3,7,11\ldots\}$. More precisely, if we now consider the cable-graph $[z,\infty)$ (for some $z<0$), $x=0$ and $y=1$, the switching lemma states that for some explicit constant $C$ (that can be shown to equal to 1):
@@
    One would think that Jim Pitman and Marc Yor would have come up instantaneously with proofs along the lines that we give in this section via explicit Laplace transform computations. As it turns out, in what probably is one of the joint famous summer papers [47], they point out exactly such a decomposition of Bessel Bridges (of general dimension), by recognizing the different terms in the Laplace transform! In particular, Formula (1.f) in the case $d=1$ gives the parity lemma for the cable graph $[0,1]$. When one conditions a Brownian bridge not to hit the origin, one gets exactly a bridge of a three-dimensional Bessel process, so that Formula (1.f) in the case $d=3$ for the decomposition of such a Bessel bridge gives exactly the above corollary; for both cases, one just has to remember the formula for the Gamma function evaluated at half-integer points to read in the Formula (1.f) of [47] that the number of excursions joining 00 and $1$ is a Poisson random variable respectively conditioned to be even or odd. It is interesting and nice to see how Marc Yor’s favourite techniques can be also reinterpreted via loop-soup decompositions. This is the type of “quest for mysteriously hidden pathwise explanations” of identities in law between functionals of stochastic processes that he was fond of.
    Another avenue to proving the general switching property could actually be to use this case of the graph with one edge and ideas from [41] to use “equivalence between effective resitance” (this is also related to the arguments in [1]), but all these different routes are of course interrelated (since the Pitman-Yor computations bear similarities with the ones presented in the previous section).
@@ 5. Outlook: An explicit bijection via “peeling”
This question is reminiscent of papers by Warren, Yor, Aldous and others [60, 61, 5] that were studying the conditional distribution of one Brownian motion (or excursion) up to some given (stopping) time $T$ given its local time profile at $T$ – one idea being to try to progressively draw part of the Brownian motion, that then eats up progressively the allotted local time, and to describe the joint dynamics of the pair (Brownian motion, remaining allotted local time). Here, the question is how to do something similar for collections of Brownian loops on graphs. This goes (initially) very much along the same lines as [2], see also [1].
@@
The obtained trajectory is then the “reversed vertex reinforced jump-process” as introduced by Sabot and Tarrès in [53] and further studied in [38]. In the limit where one lets $K\to\infty$, it is explained by Lupu, Sabot and Tarrès [39] that the process converges to a particular continuous self-interacting process. As opposed to the more strongly “locally self-repelling” processes such as the one defined and studied in [59], the sample trajectories of this process will look very much like Brownian ones (indeed, the “annealed version” will just give rise to the Brownian excursions away from $x$ by the Brownian loops, so that they always will have a finite quadratic variation), but it can be viewed as a Brownian motion with a rather complicated drift that is determined by its “remaining local time function” (the initial local time profile minus the local time of the past of the process). This updated remain profile will decrease when the process goes along. This process is well-defined and the convergence to this continuous process holds true up to the possible first time at which the local time at the current position of the process hits 00. Note that after that time, the process can not come back to this point anymore (as it would otherwise create a positive local time at this point, which is not possible), so that it has “to decide” in which direction to continue. Since this decision is “locally made” based on the local time profile, it is quite natural to guess that it will choose one of the two directions with equal probability.
@@ 6. Some first consequences for $n$-point functions and other features
One general remark is that up to the present paper, the study of large sign clusters in ${\mathbb{Z}}^{d}$ was focusing first on the derivation of connection probabilities between points, or between points and sets in the case of one-arm probabilities. For instance, a large part of the recent papers [15, 16, 17, 13, 9, 10, 11, 22] is devoted on establishing bounds for the “one-arm” estimates, i.e. for the behaviour of the probability of the event $A_{N}$ that the sign cluster containing a given point is not contained in the ball of radius $N$ around this point when $N\to\infty$). Those proofs were mostly building on Lupu’s two-point function (and in some case on an extension from [41]) on the one hand, and on insightful delicate explorations of the GFF, involving very non-trivial renormalization arguments. With the multiple-point function estimates in hand (that they then subsequently obtain), it is then possible (see in particular very recent papers [17, 11, 22]) to extract information about the large-scale properties of “typical” large clusters. This last step (from $n$-point functions to the properties of clusters) is not really specific to this loop-percolation case (see for instance the considerations of [4] for Bernoulli percolation).
@@ 6.3. The continuum versions
We conclude with the following point: In dimensions $d=3,4,5$, one expects (see [63]) the collection of loop-soup clusters to have a scaling limit when the mesh of the lattice goes to 00 (just as in $d=2$, see [37]), that will be coupled with a realization of the Brownian loop-soup in the continuum [and the properties of this scaling limit are then directly related with the large-scale properties of the models on cable-graphs]. In that context, given that the mass of excursions joining two far-away points in the cable-graph goes to 00 in the scaling limit, one expects that in the continuum, the switching property would hold in a similar way to Theorem 1: Conditioning two points $x$ and $y$ to be in the same “continuum sign-cluster” would amount to just adding one Brownian excursion joining these two points. As we have already briefly mentioned in the introduction, this has actually already be shown to hold in dimension 2 (and a domain with finite Green’s function) by Jégo, Lupu and Qian [25]. We plan to address such “switching property in the continuum setting results” in subsequent work [65].
@@ Acknowledgments.
This research has been supported by a Research Professorship of the Royal Society. I thank Juhan Aru, Hugo Duminil-Copin and Titus Lupu for their comments on preliminary versions of this paper. Many thanks also to Elie Aidékon for pointing out the Pitman-Yor paper [47].